What velocity would a 60 kg person need to move at in order to have the same momentum as a 600 nm photon? (to 2 s.f and in m s−1)
1
Expert's answer
2019-05-23T09:40:09-0400

Solution. Photon momentum can be found using the formula


"p=\frac {h} {\lambda}"

where h=6.63*10^(-34) J*s; lambda=600*10^(-9)m  is the photon's wavelength.

Human impulse can be found using the formula


"p=mv"

where m=60kg is mass of person; v is velocity. Therefore


"\frac {h} {\lambda}=mv \Longrightarrow v= \frac {h} {\lambda m}."

Hence


"v= \frac {6.63 \times 10^{-34} } {600 \times 10^{-9} \times 60}=1.84 \times 10^{-29} \frac {m} {s}"


Answer.


"1.84 \times 10^{-29} \frac {m} {s}"